OpenStudy (vampirediaries):

HELP!!!!! :) :( Use an inverse matrix to solve each equation or system.

OpenStudy (vampirediaries):

21. \[\left[\begin{matrix}3 & 5 \\ 6 & 2\end{matrix}\right]X=\left[\begin{matrix}-2 & 6 \\ 4 & 12\end{matrix}\right]\] 23. \[\left[\begin{matrix}4 & 1 \\ 2 & 1\end{matrix}\right]\left[\begin{matrix}x \\ y\end{matrix}\right]=\left[\begin{matrix}10 \\ 6\end{matrix}\right]\]

OpenStudy (vampirediaries):

@KayleeNeedsHelp , @nermaljean99

OpenStudy (vampirediaries):

@campbell_st , @ganeshie8 help, please